OpenStudy (yb1996):

Determine the values of "a" and "b" for which the following limit is true: limit as x approaches infinity of ((ax+1)/(bx-1))^x = 4

OpenStudy (anonymous):

\[\left(\frac{ax+1}{bx-1}\right)^x=4\]?

OpenStudy (yb1996):

yes

OpenStudy (anonymous):

find the limit in \(a\) and \(b\)

OpenStudy (yb1996):

how do you do that?

OpenStudy (anonymous):

oh this is a pain start with by taking the log to get \[x\ln\left(\frac{ax+1}{bx-1}\right)\]

OpenStudy (yb1996):

ok, and the other side would equal ln(4), right?

OpenStudy (yb1996):

wait, no it wouldn't

OpenStudy (anonymous):

yeah now i am confused

OpenStudy (anonymous):

we need to compute that limit in \(a\) and \(b\) and it is not of the form to use l'hopital because we don't know if \(a>b\) or if \(a<b\) or what

OpenStudy (anonymous):

so now i am thinking this method may not be right

OpenStudy (yb1996):

why would you have to know whether a >b or not to use l'hopital's rule?

OpenStudy (yb1996):

wouldn't l'hopital's rule work for any values of a and b?

OpenStudy (yb1996):

can anyone else help? satellite73 went offline....

OpenStudy (matlee):

I dont know any of this the most i can do is wolfram alpha

ganeshie8 (ganeshie8):

\(a=b=-\ln 4\) seems to work

OpenStudy (yb1996):

how did you get that?

ganeshie8 (ganeshie8):

that doesn't work, try this : \(a=b = \dfrac{1}{\ln 4}\)

ganeshie8 (ganeshie8):

so that doesn't work either.. wait il explain once i figure it out

OpenStudy (yb1996):

ok

OpenStudy (yb1996):

ok, is there any way using calculus to get that answer?

ganeshie8 (ganeshie8):

remember how to do long division ?

OpenStudy (yb1996):

yes

ganeshie8 (ganeshie8):

\[\dfrac{ax+1}{bx-1}=?\]

OpenStudy (yb1996):

would that be \[\frac{ a }{ b } + \frac{ 1-\frac{ a }{ b } }{ bx-1 }\] ?

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

quotient is correct but remainder doesn't look correct..

ganeshie8 (ganeshie8):

bx - 1 | ax + 1 | a/b ax - a/b ------------------ 1 + a/b (ax+1)/(bx-1) = a/b + (1 + a/b)/(bx-1) yes ?

OpenStudy (yb1996):

true, I messed up while subtracting

ganeshie8 (ganeshie8):

\[ \left(\frac{ax+1}{bx-1}\right)^x\\~\\ =\left(\frac{a}{b}+\frac{1+a/b}{bx-1}\right)^x\\~\\ \]

ganeshie8 (ganeshie8):

before proceeding further, let me ask you a question

ganeshie8 (ganeshie8):

as you make \(n\) large, what number does the below expression approach ? \[(0.9)^n\]

OpenStudy (yb1996):

zero

ganeshie8 (ganeshie8):

good. how about below expression : \[(1.1)^n\] what number does that approach as you make \(n\) large ?

OpenStudy (yb1996):

I presume that should go to infinity, right?

ganeshie8 (ganeshie8):

Excellent! as you can see, \((1\pm x)^n\) is either \(0\) or \(\infty\) for nonzero \(x\) as \(n\to\infty\)

ganeshie8 (ganeshie8):

Keeping that in mind, lets look at our problem

ganeshie8 (ganeshie8):

\[ \left(\frac{ax+1}{bx-1}\right)^x\\~\\ =\left(\frac{a}{b}+\frac{1+a/b}{bx-1}\right)^x\\~\\ \]

ganeshie8 (ganeshie8):

As \(x\to \infty\), is there any hope for above expression to be \(4\) if \(\dfrac{a}{b}\) is greater than \(1\) ?

OpenStudy (yb1996):

no, because you would be adding a number to 4, so if a/b is greater than 1, than the result will always be greater than 4

ganeshie8 (ganeshie8):

In other words, for \(\dfrac{a}{b}\gt 1\), the limit "diverges"

ganeshie8 (ganeshie8):

what happens if \(\dfrac{a}{b}\lt 1\) ?

OpenStudy (yb1996):

I'm guessing that it would approach 4?

ganeshie8 (ganeshie8):

My guess is that it would approach 0

ganeshie8 (ganeshie8):

as you can see, i want to argue that \(\dfrac{a}{b}\) has to equal to \(1\)

OpenStudy (yb1996):

yes, I understand, if you make b arbitrarily big, then the limit of the whole thing would approach zero, that makes sense

OpenStudy (yb1996):

so if the limit does not go to 4 when a/b is greater than 1 or less than 1, that means a/b must equal 1. Which means that a must equal b.

ganeshie8 (ganeshie8):

Yes, plug that in \(a=b\)

ganeshie8 (ganeshie8):

\[ \left(\frac{ax+1}{bx-1}\right)^x\\~\\ =\left(\frac{a}{b}+\frac{1+a/b}{bx-1}\right)^x\\~\\ \] \(a=b\), so : \[ \left(1+\frac{1+1}{bx-1}\right)^x \]

OpenStudy (yb1996):

ok, so that would simplify to \[(1+\frac{ 2 }{ bx-1 })^x\]

OpenStudy (yb1996):

would I have to take a common denominator next?

OpenStudy (yb1996):

actually, taking a common denominator will bring me back to the original equation

ganeshie8 (ganeshie8):

we may use this : \[\lim\limits_{x\to\infty} (1+\dfrac{t}{x})^x = e^t\]

ganeshie8 (ganeshie8):

but it seems we need to massage a bit before using that

OpenStudy (yb1996):

ok, and how exactly would we denominator to equal to exponent?

ganeshie8 (ganeshie8):

\[ \left(\frac{ax+1}{bx-1}\right)^x\\~\\ =\left(\frac{a}{b}+\frac{1+a/b}{bx-1}\right)^x\\~\\ \] \(a=b\), so : \[ =\left(1+\frac{1+1}{bx-1}\right)^x\\~\\ =\left(1+\frac{2/b}{x-1/b}\right)^x\\~\\ \]

ganeshie8 (ganeshie8):

fine with above ?

OpenStudy (yb1996):

yeah, you just divided both top and bottom with b

OpenStudy (freckles):

way I went about it: \[\lim_{x \rightarrow \infty}(\frac{ax+1}{bx-1})^x=4 \\ \text{ Take }\ln( ) \text{ of both sides: } \\ \ln[\lim_{x \rightarrow \infty}(\frac{ax+1}{bx-1})^x]=\ln(4 ) \\ \text{ you can bring the limit outside } \log \\ \lim_{x \rightarrow \infty} \ln[(\frac{ax+1}{bx-1})^x]=\ln(4) \\ \text{ use power rule for } \log \\ \lim_{x \rightarrow \infty} x \ln (\frac{ax+1}{bx-1})=\ln(4) \\ \lim_{x \rightarrow \infty} \frac{\ln(\frac{ax+1}{bx-1})}{\frac{1}{x}} =\ln(4) \\ \text{ In order for the limit \to actually exist we need } \frac{ax+1}{bx-1} \rightarrow 1 \text{ as } x \rightarrow \infty \\ \text{ so we can have } \frac{0}{0} \\ \text{ and use l'hospital rule } \\ \text{ so this does give you } \frac{a}{b}=1 \implies a=b \\ \text{ So you have } \\ \lim_{x \rightarrow \infty}\frac{\ln(\frac{ax+1}{ax-1})}{\frac{1}{x}}=\ln(4)\] see if you can finish apply l'hospital

ganeshie8 (ganeshie8):

\[ \left(\frac{ax+1}{bx-1}\right)^x\\~\\ =\left(\frac{a}{b}+\frac{1+a/b}{bx-1}\right)^x\\~\\ \] \(a=b\), so : \[ =\left(1+\frac{1+1}{bx-1}\right)^x\\~\\ =\left(1+\frac{2/b}{x-1/b}\right)^x\\~\\ =\left(1+\frac{2/b}{x-1/b}\right)^{x-1/b}*\left(1+\frac{2/b}{x-1/b}\right)^{1/b} \\~\\ \] taking limit gives you \[ = e^{2/b}*1 \]

OpenStudy (yb1996):

Thank you freckles, I'll try that as soon as I'm done with ganeshie8's way

OpenStudy (yb1996):

Ok, so I'm a bit confused as to how the second term equals 1

OpenStudy (yb1996):

would it be because as x is approaching infinity (2/b)/(x-1/b) is approaching zero, so 1^(1/b) would just be 1?

ganeshie8 (ganeshie8):

Yes. notice that 1/b is a constant with respect to x : \[\lim\limits_{x\to\infty}\left(1+\frac{2/b}{x-1/b}\right)^{1/b} = \left(\lim\limits_{x\to\infty}1+\frac{2/b}{x-1/b}\right)^{1/b}\\~\\~\\ = (1+0)^{1/b}\\~\\ =1 \]

OpenStudy (yb1996):

ok, that makes sense

ganeshie8 (ganeshie8):

recall, for continuous functions, we have : \[\lim f(g(x)) = f(\lim g(x))\]

ganeshie8 (ganeshie8):

\[\lim [f(x)]^n = [\lim f(x)]^n\]

OpenStudy (yb1996):

ok so lim[e^(2/b)] equals [lim e]^(2/b) ?

ganeshie8 (ganeshie8):

No, we're done. the limit evaluates to \(e^{2/b}\), simply set this equal to \(4\) and solve \(b\)

OpenStudy (yb1996):

i got b = 2/ ln(4)

ganeshie8 (ganeshie8):

Looks good !

OpenStudy (yb1996):

so how does that simplify to 1/ln(2) ?

ganeshie8 (ganeshie8):

use below log property : \[\ln(x^a) = a*\ln(x)\]

OpenStudy (yb1996):

ok, so it looks like it should be 1/ln(4^(1/2)), but I'm not sure whether that would be the correct way to do it

ganeshie8 (ganeshie8):

Easy... ln(4) = ln(2^2) = 2ln(2)

ganeshie8 (ganeshie8):

because 4 = 2^2

OpenStudy (yb1996):

OHHHH, lol, that was literaly the stupidest question I have asked.....

OpenStudy (yb1996):

Thank you so much for your help and your time!

ganeshie8 (ganeshie8):

haha thats okay :) my stupidiest question was way stupidier than ur stupidiest q

ganeshie8 (ganeshie8):

do scroll up and see freckles method using Lhopital rule...

OpenStudy (yb1996):

yes, I'm actually looking at his method right now